Progress Toward the Milestones of the enclosed schedule:
NOTE: Meetings are held every Tuesday in CEBAF Center,
Room #A110, at 11 a.m., unless otherwise notified.
NOTE: A reminder to all principals to submit a brief weekly progress to D. Gilchrist each Friday by noon
ATTENDEES: M. Augustine, G. Biallas, S. Benson, C. Bohn,
L. Dillon-Townes, B. Dunham, D. Engwall, K. Hovater, K. Jordan,
G. Krafft, R. Legg, G. Neil, C. Piller, M. Shinn,
T. Siggins, C. Sinclair, J. Song, R. Walker, M. Wiseman
The official policy on voltage applied to the gun is that the initial run Transverse experiment will be run at 250 kV after processing at 300 kV. At the end of the run, and when additional ceramics are in hand, an attempt will be made to run at 350 kV. After beam line reconfiguration, the Longitudinal experiment will be run at 350 kV after conditioning at 400 kV. If we can combine the experiments, additional Transverse Data will be taken at 350 kV and if additional ceramics are on hand at the end of the run, an attempt will be made to gather both sets of data at 500 kV.
The higher voltage, 9 MV/m cavity will be the first cavity in the quarter cryomodule. This configuration conforms to the simulations.
Safety Action Items:
New Safety Action Items:
M. Shinn should attend the Safety Warden training on August
13 to get ready to be the Safety Warden form September through
Completed Safety Action Items:
T. Siggins/D. Engwall - Get the gun and line baked and get the Experiment running. The bake was successful. The gun started cooling on Monday while the beam line remains at high temperature while its pressure continues to fall. For some work , D. Engwall will work R. Legg to get work done in series to get the gun operational faster. T. Siggins will utilize personnel from outside the group if necessary to remove the bake apparatus and will notify Rich Hall when the diagnostics can be brought back to operational status.
G. Biallas - Make a new ceramic stack and make a new cathode support tube with a anti field emission coating. Contract for ion implantation work at Berkeley is in process and samples sent to them. T. Siggins has an estimate of $2000 per ceramic to braze the Kovar rings on. The apparatus to rotate the ceramic during ion implantation will be done by us and may become a critical path item. We will use the existing stainless steel support tube material because polishing specification is not expected to be at our former 1 µm diamond polish level of perfection. Polishing may be able to be done by Electron beam polishing, a technique shown at the Voltage in a Vacuum Conference. T. Venhaus will do a seminar on the conference when he is ready.
R. Legg - Investigate an alternate scheme to use the Longitudinal apparatus to perform the high voltage portions of the Transverse experiment. B. Dillon Townes will put together a drawing of the proposed line , working from information from D. Kehne. D. Kehne will get the proposal done by Aug. 22.
Determine the need for additional personnel required to match or beat the Schedule. No needs this week.
T. Siggins - Write the tricks of polishing learned at SLAC. No action.
G. Biallas - Finish the LCW System: Flow & Temp reading, Level indicators, Full qualification walk through needed. No action.
K. Jordan/S. Benson - Failure analysis for the 10 MeV beam handling. No action
M. Shinn - Put phase noise documentation on the web. In process.
All - general cleanup after bake. Checkout systems, straighten out cabling. Partially done.
R. Legg - Publish the straw man for the 10 MeV Experiment . Files need to be compatible with the Web.
G. Biallas - integrate the new experimental schedule developed by Bob Legg into the main schedule. In process.
B. Dillon Townes - Get the 10 MeV beam line signed off. Two thirds in signature today. The remaining third is given out for comment.
K. Jordan - Get the bunch length measurement contract signed. Promised for this week.
K. Bohn - Produce the design for the multislit for 10 MeV. K. Jordan will be the Engineer. The design is started. K. Jordan will get the schedule bars to G. Biallas.
New Action Items:
D. Kehne - Present the Longitudinal simulation results at 250 kV.
C. Piller - Have the copper cavity plated or painted to withstand the 250 °C bake.
D. Kehne - In the proposed, combined 500 kV proposal, utilize the its second bakable solenoid for the solenoid in front of the Chopper rather than the built-in solenoid.
K. Bohn/R. Legg - Resolve where the Quartercryomodule should be tested first, in the Cryomodule Test Cave, testing one cavity at a time for basically a window test or to go through the disruption now and test in the ITS Enclosure in the offset position, coincident with the Longitudinal experiment.
Completed Action Items:
G. Biallas - Define Nomenclature document Definitions have to get to T. Hassler. Complete.
M. Shinn - Get HeNe at low power onto cathode. Complete.
K. Jordan - Make sure that version control of settings are maintained. This is a tough job. The CEBA still has problems. No defined solution is seen. It is in K. Jordan's plans. Complete.
Individual Reports of 8/9/96:
We are holding schedule. Bake successful. Start of high voltage conditioning starts this afternoon. Coupons of alumina being ion implanted at LBL, parameterizing the process. Kovar ring brazing source in hand. Shop is starting on new cathode support tube to be coated with anti field emission coating.
L. A. Dillon-Townes:
Tested the both injector HOM filters to 50 kW. 30 Minute heat run looks OK. .
Powered up the second HPA unit.
1. Chuck Fogleman of Wesgo advised that their schedule for delivering ceramics appeared to be slipping beyond 9/3. His language was rather vague.I asked him to move the delivery date instead ahead of 9/3 and see what he could do to expedite things. He did not get back to me yesterday and I will contact him today. I have not yet raised the question of overtime fees.
2. Tim has lined up three good brazing sources:
3. A sketch of a chamber configuration for ion implanting the inside of ceramic cylinders has been faxed to LBL for their comment and approval.
4. FM Technologies is still working on a quote for coating the cathode support tube with a field emission supressant.To speed things up I asked them to go ahead and quote the Venhaus samples separately and follow with a quote for the tube. I suggested the maximum tube length we would need to cover is that region exposed to the ceramic nearest the cathode lying between 10 and 22 inches measured from the cathode end of the tube.
C. Hovater and I continued working on running the drive laser's mode-locker through the divide-by-
k40 and MO. There are problems, which we are working on resolving.
Worked on the drive to resolve SHG power problems.
W. Ibrahim just got his high-dynamic range autocorrelator running
under computer control. He is now working through some control
New drive probes for the Chopper cavity are to be built within a few weeks. The probe coupling needs to be checked and the cavity tuned before the next test.
The additional fiber optic cable for the Arc Detectors on the 50 KW system is over due.
A discussion was held again this week on the RF system software for the IR FEL. A list of parameters will be given to the software group next week.
The plumbing was completed for the Buncher cavity water skid this week. A water leak with the Buncher cavity was uncovered and repaired.
The FEL Master Oscillator, an HP8644A, was calibrated and returned
from HP this week. It was reinstalled in the ITS.
A test of 4 klystrons and 2 circulators operating at 8 KW is
being planned and the parts are being collected.
ITS Schedule for 8/13/96